# MAYA: Addressing Inconsistencies in Generative Password Guessing through a Unified Benchmark

### v2

@@ 5.8 General Benchmarking
This section presents a comprehensive benchmark of the selected generative models, leveraging prior evaluations to facilitate direct comparisons across key metrics: performance, generalizability, and generated password quality. Table IX summarizes all aggregated results, reflecting the trends observed in the preceding sections. Regarding performance metrics, computed as the weighted average percentage of guessed passwords, FLA consistently achieves the best results across all tasks, with PassGPT being its sole direct competitor. The remaining models exhibit greater task dependency: for instance, PLR-GAN leads in in-distribution scenarios, while PassFlow excels in cross-community and cross-culture settings. However, when evaluating generalizability, the overall ranking shifts. We consider three metrics: sensitivity to training set size, measured using the coefficient of variation, and performance loss in cross-community and cross-culture scenarios, quantified as weighted average loss relative to the in-distribution setting. PassFlow consistently emerges as the most robust model across generalization tasks. Surprisingly, FLA and PassGPT exhibit the highest loss in the cross-cultural setting, with PassGPT additionally demonstrating the lowest robustness in the cross-community scenario, whereas FLA maintains an average value. Finally, concerning the quality metrics, we analyzed both uniqueness and humanness of generated passwords. While not previously emphasized in our analysis, uniqueness is a commonly adopted metric for evaluating the diversity of generated passwords. While FLA’s generation approach guarantees 100% uniqueness by design, PassFlow, leveraging the GS technique, and VGPT2 produce around 90% unique passwords. PLR-GAN and PassGPT achieve slightly above 70% uniqueness, and PassGAN demonstrates the lowest uniqueness at 54.59%. Regarding humanness distance, measured as the average percentage distance between generated and real passwords, results align with our previous observations: models such as PassGPT and PLR-GAN generate passwords closely resembling human-created ones, making them highly suitable for applications like honeywords, where it is important to deceive adversaries using realistic passwords, and as potential sources of synthetic data for future research, thereby addressing challenges with real-world password datasets (e.g., difficult to obtain, outdated, ethical concerns).
@@ 6 Insights and Lessons Learned
Sequential Architectures Perform Best
FLA and PassGPT, based on LSTM and GPT2, consistently outperform other architectures, underscoring the advantages of sequential models in capturing dependencies and generating more accurate, complex guesses.
Generative Models Surpass Traditional Tools
Early generative approaches—such as PassGAN, PassFlow, and PLR-GAN—struggles to match the performance of traditional tools. However, the emergence of transformer-based architectures has shifted the research landscape toward models capable of consistently outperforming them. PassGPT, the most recent approach evaluated in our study, generally surpasses JtR, Hashcat, PCFG, and OMEN. As research progresses, more advanced LLMs are likely to further improve password-guessing effectiveness, whereas traditional tools have likely reached their performance ceiling.
@@
Four out of six models struggle to guess complex passwords—those containing special characters or multiple character classes—and longer passwords exceeding eight characters. Only FLA and PassGPT achieve a non-negligible success rate in these cases, highlighting the persistent challenge posed by long and complex passwords and revealing model limitations.
Rare Does Not Mean Hard to Guess
@@
Generative models effectively learn to generate passwords that closely resemble human-created ones, making them valuable for a variety of tasks beyond just password-guessing.
@@ Appendix A Models and Tools Implementation
This appendix provides additional implementation details on all models. Each generative model was ported to PyTorch 2.6.0, standardizing dependencies and ensuring reproducibility. All training parameters used were as specified by the original papers/implementation, unless otherwise stated.
@@
We implemented FLA following the description of the large model mentioned in [40]. Specifically, the architecture consists of three LSTM layers with 1000 cells each, followed by two FC layers. Each checkpoint was trained for 20 epochs, processing input backwards. Unlike other generative approaches, FLA requires a probability threshold as input, filtering out passwords whose overall probability falls below it. We set this threshold to $10^{-8}$ for up to $10^{6}$ generated passwords, $10^{-9}$ for $10^{7}$, and $10^{-10}$ for $5\times 10^{8}$. After generation, we sort guesses by probability in descending order and select the first $n$ ones, equivalent to finding the optimal threshold to generate exactly $n$ passwords.
@@
We implemented PassGAN as described in the original work [29]: using an IWGAN [20] with both the generator and discriminator composed of 5 residual blocks [26]. Following the original setup, we trained PassGAN’s models for $200,000$ iterations, evaluating checkpoints every $10,000$ steps.
@@
We implemented VGPT2 [7] without introducing any architectural or training modifications.
@@
We used the publicly available implementation of OMEN [15] from [34]. We configured it as a 4-gram Markov model with a coverage set to 0, thereby relying exclusively on guesses from the Markov attack.
@@
We used the publicly available implementation of PCFG from [34]. We set the coverage parameter to 1, effectively relying solely on guesses produced by the PCFG model.
TABLE XII: Marginal gain in successful matches as the number of generated passwords increases, with each column representing generation intervals (e.g., 1M-2.5M indicates the marginal gain when expanding from 1M to 2.5M guesses). For a given interval X-Y, Total indicates the marginal gain achieved when increasing the number of guesses from X to Y, expressed as a percentage of the total number of test passwords. Relative indicates the marginal gain relative to the number of matches obtained at X. Results are averaged across all password lengths.
Model 1M-2.5M 2.5M-5M 5M-7.5M 7.5M-10M 10M-25M 25M-50M 50M-75M 75M-100M 100M-250M 250M-500M Total Relative Total Relative Total Relative Total Relative Total Relative Total Relative Total Relative Total Relative Total Relative Total Relative PassGAN +0.49 +83.75 +0.55 +50.94 +0.40 +24.84 +0.32 +15.95 +1.23 +53.74 +1.14 +32.85 +0.75 +16.47 +0.56 +10.69 +1.96 +34.15 +1.63 +21.38 PLR-GAN +0.62 +97.50 +0.74 +59.45 +0.57 +28.94 +0.48 +18.91 +1.94 +65.41 +1.96 +40.34 +1.34 +19.54 +1.03 +12.55 +3.67 +40.28 +3.21 +25.59 PassFlow +0.62 +69.67 +0.73 +48.88 +0.56 +25.15 +0.44 +15.72 +1.72 +52.72 +1.58 +31.53 +1.05 +15.85 +0.80 +10.38 +2.84 +33.66 +2.48 +21.97 PassGPT +1.20 +114.72 +1.60 +71.91 +1.34 +35.32 +1.18 +22.97 +5.21 +83.58 +5.45 +48.49 +3.66 +22.15 +2.73 +13.56 +8.89 +39.13 +6.53 +20.63 VGPT2 +0.18 +137.88 +0.28 +90.08 +0.25 +43.08 +0.24 +28.34 +1.18 +110.51 +1.45 +65.38 +1.11 +30.52 +0.92 +19.44 +3.65 +65.02 +3.43 +37.03 FLA +4.33 +58.54 +4.17 +35.55 +2.66 +16.71 +1.93 +10.40 +6.57 +32.14 +5.55 +20.62 +3.33 +10.28 +2.34 +6.56 +7.44 +19.62 +5.84 +13.00
@@ B.1 Additional Details On RQ6
To address RQ6, we employed two metrics: the Jaccard Index and the Mergeability Index. We now present their mathematical definitions:
@@ B.2 Additional Details On RQ7
Let $D^{train}=\bigcup_{j=1}^{|D|}D^{train}_{j}$ and $D^{test}=\bigcup_{j=1}^{|D|}D^{test}_{j}$ denote the union of all training and testing datasets, respectively. Let $R$ be the set of randomly generated passwords, each between 6 and 12 characters in length. The lower-bound baseline $L$ and upper-bound baseline $U$ for a given metric $dist$ are computed as follows:
@@
Next, for each model $m_{i}$, let $P_{i}$ be the set of passwords generated by model $m_{i}$ across all datasets. The value $d_{i}$ for a model $m_{i}$ on a given metric $dist$ is then computed using the following equation:
@@
We now provide further details regarding the four selected metrics, followed by an additional analysis on IMD.
@@
IMD, which stands for Intrinsic Multi-scale Distance, is a metric designed to compare the data manifolds of two distributions. IMD provides an intrinsic method to lower-bound the spectral Gromov-Wasserstein distance between two manifolds. Unlike other approaches that focus on extrinsic properties and are uni-scale, IMD is intrinsic, meaning it is not dependent on the transformation of the manifold and multi-scale, capturing both local and global properties.
@@
MTopDiv introduced a method that tracks multiscale topology discrepancies between two distributions, P and Q, in a high-dimensional space. Unlike IMD, MTopDiv also considers extrinsic properties, using position and translation to capture structural differences between the two distributions.
TABLE XIII: IMD outputs as the distributions $P$ and $Q$ vary.
@@
To better understand the rationale behind the high values obtained with IMD across several models, we repeat the metric analysis using $10^{7}$ generated passwords. The results, shown in Table XIV, differ significantly from those obtained with $5\times 10^{8}$ generated passwords (Table VIII), particularly for PassFlow, PLR-GAN, and FLA, which exhibit notably higher IMD scores in the smaller sample. To investigate this discrepancy, we compare the length distributions of the $10^{7}$ and $5\times 10^{8}$ generated samples. We observe that, in the smaller set, models such as PassFlow and PLR-GAN—based on GS and DPG techniques, respectively—exhibit a strong bias toward shorter passwords. This behavior stems from the tendency of such models to initially saturate the space of simpler (i.e., shorter) passwords before extending to more complex ones. A similar pattern is observed with FLA, which outputs the most probable passwords. Consequently, when fewer samples are generated, the length distribution exhibits a strong bias toward shorter lengths, as shorter passwords are more probable. These findings confirm our hypothesis that IMD is highly sensitive to length distribution differences: the greater the deviation of the generated passwords’ length distribution from the real one, the higher the distance measured by IMD.
@@
We now focus on the 000webhost dataset, which yields a particularly high IMD score for PassFlow when evaluating $10^{7}$ generated passwords. Although this score exceeds that obtained from random passwords, it does not necessarily indicate that PassFlow’s outputs are random or even close to random. To demonstrate this, we compute the IMD metric between PassFlow’s generated passwords and random passwords. The results, shown in Table XIII, are compared against the two baselines and reveal that PassFlow’s passwords are farther from random passwords than from real ones. We further examined the effect of aligning the length distribution of the random passwords with that of PassFlow’s 000webhost-generated passwords. As reported in Table XIII, this adjustment results in a higher IMD score, thereby confirming that IMD is sensitive to differences in length distribution.
@@
Lastly, we investigate the impact of various degrees of mode failure, such as mode dropping and mode invention, on the IMD metric. For this analysis, we used the 19 patterns listed in Table III. As shown in Figure 10, the pattern distribution of random passwords deviates substantially from the others. Mode invention is observed in patterns r7, r9, r11, r13, r16, and r17, while mode dropping occurs in commonly observed patterns such as r6, r10, and r19. These findings suggest that the IMD metric does not adequately capture the varying degrees of mode failure and further support the hypothesis that IMD primarily outputs high values due to a model’s inability to replicate the real password length distribution. Interestingly, PassGPT performs particularly well in approximating the pattern distribution of 000webhost, with its pattern distribution closely aligning with that of the real passwords. Additionally, FLA outperforms PassFlow in capturing this distribution, with the latter struggling to replicate patterns r10, r14, and r19. These observations are consistent with the results obtained using the MTopDiv metric in Table VIII, which explicitly accounts for such mode-related discrepancies.
@@
TABLE XIV: Distance between human- and generative model-created passwords when generating $10^{7}$ passwords.

### v4

@@ 5.9 General Benchmarking
This section presents a comprehensive benchmark of the selected generative models, leveraging prior evaluations to facilitate direct comparisons across key metrics: performance, generalizability, and generated password quality. Table XI summarizes all aggregated results, reflecting the trends observed in the preceding sections. Regarding performance metrics, computed as the weighted average percentage of guessed passwords, FLA generally achieves the best results across tasks, with PassGPT, PCFG, and OMEN, as its main competitors. The remaining models exhibit greater task dependency: for instance, PLR-GAN leads in in-distribution scenarios, while PassFlow excels in cross-community and cross-culture settings. However, when evaluating generalizability, the overall ranking shifts. We consider three metrics: sensitivity to training set size, measured using the coefficient of variation, and performance loss in cross-community and cross-culture scenarios, quantified as weighted average loss relative to the in-distribution setting. PassFlow consistently emerges as the most robust model across generalization tasks. Surprisingly, FLA, PCFG, OMEN, and PassGPT exhibit the highest loss in the cross-cultural setting, with PassGPT , OMEN, and PCFG additionally demonstrating the lowest robustness in the cross-community scenario, whereas FLA maintains an average value. Finally, concerning the quality metrics, we analyzed both uniqueness and humanness of generated passwords. While not previously emphasized in our analysis, uniqueness is a commonly adopted metric for evaluating the diversity of generated passwords. While FLA’s generation approach guarantees 100% uniqueness by design, PassFlow, leveraging the GS technique, and VGPT2 produce around 90% unique passwords. PLR-GAN and PassGPT achieve slightly above 70% uniqueness, and PassGAN demonstrates the lowest uniqueness at 54.59%. Regarding humanness distance, measured as the average percentage distance between generated and real passwords, results align with our previous observations: models such as PassGPT and PLR-GAN generate passwords closely resembling human-created ones, making them highly suitable for applications like honeywords [36], where it is important to deceive adversaries using realistic passwords, and as potential sources of synthetic data for future research, thereby addressing challenges with real-world password datasets (e.g., difficult to obtain, outdated, ethical concerns).
@@ 6 Insights and Lessons Learned
DL-based Autoregressive Models Perform Best
FLA and PassGPT, based on LSTM and GPT2, consistently outperform other generative approaches, underscoring the advantages of deep learning autoregressive models in capturing dependencies and generating more accurate, complex guesses.
Generative Approaches Surpass Traditional Rule-Based Tools
Early generative approaches—such as PassGAN, PassFlow, and PLR-GAN— struggle to match the performance of traditional rule-based tools. However, the emergence of transformer-based architectures has shifted the research landscape toward models capable of consistently outperforming them. PassGPT, the most recent approach evaluated in our study, generally surpasses JtR and Hashcat. As research progresses, more advanced LLMs are likely to further improve password-guessing effectiveness, whereas traditional tools have likely reached their performance ceiling.
Traditional ML-Based Approaches Remain Competitive
Traditional ML-based approaches remain widely competitive, with PCFG and OMEN generally ranking between 2nd and 4th across each benchmarked metric. Furthermore, they also demonstrate strong generalizability and, for PCFG, good performance on longer passwords and complex patterns.
Rare Does Not Mean Hard to Guess
@@
Four out of eigth generative models struggle to guess complex passwords—those containing special characters or multiple character classes—and longer passwords exceeding eight characters. Only FLA, PassGPT, and the ML-based tools OMEN and PCFG achieve a non-negligible success rate in these cases, highlighting the persistent challenge posed by long and complex passwords and revealing model limitations.
Models Reach a Hard Ceiling Even with Larger Datasets
Based on the insights obtained in RQ3, we expect larger datasets to further improve performance for certain models, including PassGPT, VGPT2, and PCFG. Being an advanced autoregressive model, PassGPT is expected to scale particularly well, following trends observed in NLP [37, 31]. However, all models are likely to reach a hard ceiling as they approach the limits of modeling complex/infrequent passwords that lie on the extremities of the learned distribution, which cannot be fully captured even with larger datasets.
@@
The selected models have been shown to effectively learn to generate passwords that closely resemble human-created ones, making them valuable for a variety of tasks.
@@
On Practical Implications
To improve password security, real-world systems usually rely on password policies and complementary mechanisms such as honeywords [36], password strength meters [7, 9], and password blocklists. Our RQ7 findings show that some approaches, such as PassGPT, can generate exceptionally human-like passwords. This ability is directly relevant for strengthening the generation of realistic honeywords [17] and improving existing password blocklists with easily-guessable passwords. Additionally, effective generative models can also be leveraged to improve the effectiveness of password strength meters [44].
@@ 7 Conclusion
Acknowledgments
This work was supported by (1) Project SERICS under the National Recovery and Resilience Plan by the Minister of University and Research (NRRP MUR) Program funded by the European Union-NextGenerationEU under Grant PE00000014, (2) the NRRP MUR Program funded by the European Union-NextGenerationEU, Mission 4, Component 1, CUP B53C24002200004, and (3) the "Progetto di Ricerca Grandi 2022, Sapienza University of Rome".
@@ Appendix A Models and Tools Implementation
This appendix provides additional implementation details for each selected model and presents a categorization of each approach in Table I. Regarding deep generative models, we standardized dependencies by porting them to PyTorch 2.6.0, as some were originally implemented in TensorFlow or outdated PyTorch versions. To provide a unified comparison framework, we designed an interface and adapted each model accordingly. Future research can easily integrate new models by subclassing the base model class and implementing the required abstract methods. Unless otherwise stated, all training parameters are the same as those used in the original works. We defined a validation procedure for our implementation and assessed its accuracy by comparing our results with those reported in the original papers, observing negligible variations falling within the margin of error. The sole exception was VGPT2, for which we were unable to replicate the published results using the official implementation.
@@
We implemented FLA following the description of the large model mentioned in [48]. Specifically, the architecture consists of three LSTM layers with 1000 cells each, followed by two FC layers. Each checkpoint was trained for 20 epochs, processing input backwards. Unlike other generative approaches, FLA requires a probability threshold as input, filtering out passwords whose overall probability falls below it. We set this threshold to $10^{-8}$ for up to $10^{6}$ generated passwords, $10^{-9}$ for $10^{7}$, and $10^{-10}$ for $5\times 10^{8}$. After generation, we sort guesses by probability in descending order and select the first $n$ ones, equivalent to finding the optimal threshold to generate exactly $n$ passwords. Sorting $5\times 10^{8}$ passwords is highly computationally demanding; therefore, we implemented a custom min-heap in Cython, which significantly improved FLA’s resource efficiency.
@@
We implemented PassGAN as described in the original work [33]: using an IWGAN [23] with both the generator and discriminator composed of 5 residual blocks [29]. We trained PassGAN’s models for 200,000200,\!000 iterations, validating checkpoints every 10,00010,\!000 steps.
@@
We implemented VGPT2 [8] without introducing any architectural or training modifications. We followed the official code for the implementation, unifying the testing methodology with our framework.
@@
We used the publicly available implementation of OMEN [18] from [41]. Following [18], we configured it as a 4-gram Markov model with a coverage set to 0.
@@
We used the publicly available implementation of PCFG from [41]. We set the coverage parameter to 1, thereby relying exclusively on guesses produced by the PCFG model.
@@ B.1 Additional Details On RQ6
To address RQ6, we employed two metrics: the Jaccard Index and the Mergeability Index. We now present their formal definitions:
@@ B.2 Additional Details On RQ7
Let $D^{train}=\bigcup_{j=1}^{|D|}D^{train}_{j}$ and $D^{test}=\bigcup_{j=1}^{|D|}D^{test}_{j}$ denote the union of all training and testing datasets, respectively. Let $R$ be the set of randomly generated passwords between 6 and 12 characters in length. The lower ($L$) and upper ($U$) baselines for a metric distdist are computed as follows:
@@
Next, for each model $m_{i}$, let $P_{i}$ be the set of passwords generated by model $m_{i}$ across all datasets. The value $d_{i}$ for a model $m_{i}$ on a given metric distdist is then computed using the following equation:
@@
Intrinsic Multi-scale Distance is a metric designed to compare the data manifolds of two distributions. IMD provides an intrinsic method to lower-bound the spectral Gromov-Wasserstein distance between two manifolds. Unlike other approaches that focus on extrinsic properties and are uni-scale, IMD is intrinsic, meaning it is not dependent on the transformation of the manifold and multi-scale, capturing both local and global properties.
@@
MTopDiv measures multiscale topology discrepancies between two distributions, P and Q, in a high-dimensional space. Unlike IMD, MTopDiv also considers extrinsic properties, using position and translation to capture structural differences between the distributions.
TABLE XV: IMD outputs as the distributions $P$ and $Q$ vary.
@@
To better understand the rationale behind the high values obtained with IMD across several models, we repeat the metric analysis using $10^{7}$ generated passwords. The results, shown in Table XVI, differ significantly from those obtained with $5\times 10^{8}$ generated passwords (Table X), particularly for PassFlow, PLR-GAN, and FLA, which exhibit notably higher IMD scores in the smaller sample. To investigate this discrepancy, we compare the length distributions of the $10^{7}$ and $5\times 10^{8}$ generated samples. We observe that, in the smaller set, models such as PassFlow and PLR-GAN—based on GS and DPG techniques, respectively—exhibit a strong bias toward shorter passwords. This behavior stems from the tendency of such models to initially saturate the space of simpler (i.e., shorter) passwords before extending to more complex ones. A similar pattern is observed with FLA, which outputs the most probable passwords. Consequently, when fewer samples are generated, the length distribution exhibits a strong bias toward shorter lengths, as shorter passwords are more probable. These findings confirm our hypothesis that IMD is highly sensitive to length distribution differences: the greater the deviation of the generated passwords’ length distribution from the real one, the higher the distance measured by IMD.
@@
We now focus on the 000webhost dataset, which yields a particularly high IMD score for PassFlow when evaluating $10^{7}$ generated passwords. Although this score exceeds that obtained from random passwords, it does not necessarily indicate that PassFlow’s outputs are random or even close to random. To demonstrate this, we compute the IMD metric between PassFlow’s generated passwords and random passwords. The results, shown in Table XV, are compared against the two baselines and reveal that PassFlow’s passwords are farther from random passwords than from real ones. We further examined the effect of aligning the length distribution of the random passwords with that of PassFlow’s 000webhost-generated passwords. As reported in Table XV, this adjustment results in a higher IMD score, thereby confirming that IMD is sensitive to differences in length distribution.
@@
Lastly, we investigate the impact of various degrees of mode failure, such as mode dropping and mode invention, on the IMD metric. For this analysis, we used the 19 patterns listed in Table IV. As shown in Figure 10, the pattern distribution of random passwords deviates substantially from the others. Mode invention is observed in patterns r7, r9, r11, r13, r16, and r17, while mode dropping occurs in commonly observed patterns such as r6, r10, and r19. These findings suggest that the IMD metric does not adequately capture the varying degrees of mode failure and further support the hypothesis that IMD primarily outputs high values due to a model’s inability to replicate the real password length distribution. Interestingly, PassGPT performs particularly well in approximating the pattern distribution of 000webhost, with its pattern distribution closely aligning with that of the real passwords. Additionally, FLA outperforms PassFlow in capturing this distribution, with the latter struggling to replicate patterns r10, r14, and r19. These observations are consistent with the results obtained using the MTopDiv metric in Table X, which explicitly accounts for such mode-related discrepancies.
@@
TABLE XVI: Distance between human- and generative model-created passwords when generating $10^{7}$ passwords.